OpenStudy (anonymous):

solve this -5-4x+7x=1

OpenStudy (anonymous):

-5+3x=1 3x=1+5 3x=6 x=6/3 x=2

OpenStudy (anonymous):

how did u get that? do i have to add 4x and 7x?

OpenStudy (anonymous):

You have -4x and +7x.

OpenStudy (anonymous):

So if you have 7 x's and you take 4 away, how many x's do you have?

OpenStudy (anonymous):

yeah or is it suppose to be sqrd?

OpenStudy (anonymous):

its not negative 4! it is positive 4x

OpenStudy (anonymous):

-5-4x+7x = -5 + (-4x) + 7x

OpenStudy (anonymous):

negative 5 minus 4x plus 7x equals 1

OpenStudy (anonymous):

Yes, but subtracting 4x is the same as adding negative 4 x.

OpenStudy (anonymous):

so everytime you have a number and a variable, u have to subtract it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

If I say -5 - 4 + 7 isn't that the same as -5 + -4 + 7?

OpenStudy (anonymous):

yea! !

OpenStudy (anonymous):

So -5 - 4x + 7x = -5 + (-4x) + 7x

OpenStudy (anonymous):

-5-4x+7x=1 -5-11x=1 -5 -5 11x=-4 4 4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

show me how u did it

OpenStudy (anonymous):

7x + -4x is not -11x

OpenStudy (anonymous):

its not negative 11! its positive 11x!

OpenStudy (anonymous):

You are fundamentally not understanding addition and subtraction with regards to positive and negative numbers. Please review this material either in your book, or via the khanacademy.org videos.

OpenStudy (anonymous):

x=2